# Causal machine learning for heterogeneous treatment effects in the presence of missing outcome data

### v1

@@ 3 DR-learner/EP-learner extensions
In this section, we extend the DR-learner (Kennedy, 2023) and EP-learner (van der Laan et al., 2024) to handle missing outcome data, leading to two new estimators, the mDR-learner (missing outcome DR-learner) and mEP-learner (missing outcome EP-learner) respectively.
@@
We begin by considering an extension of the DR-learner to the MAR outcome data setting. Recall that the DR-learner offers a one-step EIF based estimator of the CATE, constructing pseudo-outcomes using the EIF of the MSE for the CATE, which it regresses on the variables defining the CATE. By doing so, the DR-learner minimizes the MSE risk function and offers reduced sensitivity to errors in its nuisance functions, allowing data adaptive techniques to be used. This said, when outcomes are MAR, the DR-learner fit using the available case analysis does not account for the under-representation which occurs as a result of missing outcomes. This limitation occurs as the risk which the DR-learner uses to derive its pseudo-outcomes takes a new form when outcomes are MAR,
$\psi_{CATE}=\operatorname{\mathbb{E}}\left[\theta(X)^{2}-2\theta(X)(\operatorname{\mathbb{E}}[Y|C=1,A=1,Z]-\operatorname{\mathbb{E}}[Y|C=1,A=0,Z])\right],$ (8)
now involving the indicator for outcomes being non-missing. Consequently, the EIF of this risk function in this setting does not correspond to the EIF in the complete data setting. In Appendix D we derive the form of this EIF, showing it to be:
@@
Using this EIF, we can construct pseudo-outcomes for the mDR-learner which account for groups which are under-represented as a result of missing outcome data:
@@
We note that IPCWs now appear in combination with IPTWs, allowing these pseudo-outcomes to account for the shift in covariate distribution caused by both missing outcome data and confounding. The mDR-learner then proceeds by regressing the estimates of these pseudo-outcomes $Y_{mDR}$ against covariates $X$ to obtain estimates of the CATE (See Algorithm 1).
@@
For the mDR-learner to obtain oracle efficiency, it requires that the product of the errors from its outcome models and propensity score model, as well as the product of the errors from its outcome models and missingness model, converge at a faster rate than oracle estimates (i.e. estimates obtained from the second stage pseudo-outcome model under known nuisance functions). For instance, consider an example where the oracle CATE estimates converge at a $\sqrt{n}$ rate. In this example, if we wish to estimate the nuisance functions within the mDR-learner without impacting the overall convergence rate of our CATE estimates, then we require the estimators for each of our nuisance functions to converge at rates faster than $4]{n}$, hence allowing the product of errors to converge at a rate faster than $\sqrt{n}$. Equally, when the outcome function estimators converge at rates slower than this, oracle efficiency can still be obtained as long as the propensity score estimator and missingness function estimator converge at rates which are sufficiently fast for the above conditions to hold. Under these weakened convergence requirements, ML can then be used when estimating nuisance functions within the mDR-learner without errors propagating through to the CATE.
Algorithm 1 mDR-learner algorithm
1:Split the data randomly into $K$ (e.g., 10) equal sized folds of $n$ observations from $O=(Z,A,YC)$, denoted $D_{1},\ldots,D_{K}$.
[…]
4:Construct the pseudo outcomes for each individual in the data using
$Y_{mDR}=\frac{\left(A-\hat{\pi}(Z)\right)C}{\hat{\pi}(Z)\left(1-\hat{\pi}(Z)\right)\hat{G}(A,Z)}\left\{Y-\hat{\mu}^{A}(Z)\right\}+\hat{\mu}^{1}(Z)-\hat{\mu}^{0}(Z)$
5:Regress the pseudo outcomes $Y_{mDR}$ on covariates $X$, and obtain predictions of $\theta(X)$:
$\hat{\theta}_{mDR}(X)=\hat{\operatorname{\mathbb{E}}}[Y_{mDR}|X]$
@@
We now demonstrate how the EP-learner, a targeted-learning based framework can be extended for the setting with MAR outcome data. As outlined in Section 2.2, this framework uses a two-step approach, first defining an iTMLE procedure to remove plug-in bias, then regressing the pseudo-outcomes obtained from this procedure against $X$ to obtain CATE estimates. However, when outcomes are MAR, we encounter the same problem discussed for the DR-learner, which is that the EIF used to derive its pseudo-outcomes now takes a new form, eq. (9). This EIF contains IPCWs within its drift term (second term), and hence, if we wish to set the sample average of the drift term to 0, removing plug-in bias, we must update the iTMLE process used within the EP-learner.
To update the iTMLE process, we re-define the weight using eq. (9): $\hat{H}(A,C,Z)=\frac{CA}{\hat{G}(A,Z)\hat{\pi}(Z)}+\frac{C(1-A)}{\hat{G}(A,Z)\left(1-\hat{\pi}(Z)\right)}$. By using this weight within the same iTMLE algorithm outlined previously, we ensure that the sample average of the second term in eq. (9) converges 0, which in turn means that the mEP-learner will be equivalent to the oracle efficient one-step mDR-learner when its nuisance function estimates converge sufficiently fast. The mEP-learner algorithm is provided in Algorithm 2, and for a general proof of this equivalence for EP risks, see the original EP-learner paper (van der Laan et al., 2024).
Algorithm 2 mEP-learner algorithm
1:Split the data randomly into $K$ (e.g., 10) equal sized folds of $n$ observations from $O=(Z,A,YC)$, denoted $D_{1},\ldots,D_{K}$.
2:For $j\in 1,...,K$ and using all folds $\{D_{i};i=1,\ldots,K,i\neq j\}$ except $D_{j}$, train models for
[…]
4:For all individuals in the data, update the outcome predictions
  1. a)
    Calculate the clever covariate, $\hat{H}(A,C,Z)=\frac{CA}{\hat{G}(A,Z)\hat{\pi}(Z)}+\frac{C(1-A)}{\hat{G}(A,Z)\left(1-\hat{\pi}(Z)\right)}$.
  2. b)
    Choose a sieve basis, $\hat{\varphi}(X)$
  3. c)
    Run a linear regression of outcomes $Y$ on feature vector $\hat{\varphi}(X)$ with offset $\hat{\mu}^{A}(Z)$ and weight $\hat{H}(A,C,Z)$ in the complete cases.
  4. d)
5:Construct the pseudo outcome for all individuals in the data
$Y_{mEP}=\hat{\mu}^{1*}(Z)-\hat{\mu}^{0*}(Z)$ (19)
6:Regress the pseudo outcomes $Y_{mEP}$ on covariates $X$, and obtain predictions of $\theta(X)$:
$\hat{\theta}_{mEP}(X)=\hat{\operatorname{\mathbb{E}}}[Y_{mEP}|X]$
@@ 3.3 Implementation
As both the mDR-learner and mEP-learner serve as general frameworks for estimating the CATE, their implementation requires the user to make several key decisions. In this section, we break these decisions down into two groups; decisions required for obtaining CATE estimates, and decisions required for assessing CATE performance.
@@ 3.3.1 Decisions required for obtaining CATE estimates
To obtain CATE estimates from either the mDR-learner or mEP-learner, the user must first decide which estimation technique they will use to estimate the nuisance functions and second stage model. As both learners are model-agnostic, this choice is flexible, and data adaptive techniques can be used. However if oracle performance is to be achieved, the estimators from the nuisance models must converge sufficiently fast to the truth. For this reason, we illustrate their implementation using the Super Learner (Van der Laan et al., 2007), an ensemble learner that allows a range of data-adaptive algorithms to be implemented and which performs asymptotically as well as its best candidate learner.
@@
After algorithm choice, users are also required to decide which type of sample splitting they will implement within the mDR or mEP learners, ensuring the data used in the pseudo-outcome regression is i.i.d. Various options exist for achieving this, including K-fold cross-validation (see Section 2.2) or independent sample splitting, which splits the data in two before training the nuisance models/second stage model on each half separately (Kennedy, 2023). However, estimators that allocate fully independent data for each nuisance/target model are typically less efficient, with the smaller sample sizes introducing finite sample bias into the estimation process. For this reason, we demonstrate the application of the mDR-learner and mEP-learner by using a 10-fold cross-validation process (see Algorithm 1), however, we note that the appropriate number of folds will depend on the complexity and smoothness of the underlying nuisance/target functions. If users wish to explore variations of cross-fitting approaches, sensitivity analyses could be run using alternative numbers of splits.
@@
We also highlight that when cross-validation is used, the reduction in training sample size for each model can introduce positivity violations, leaving certain subgroups poorly represented within some folds. To overcome this, the mDR-learner and mEP-learner can be run a large number of times ($J$ times), using a different seed for the cross-fitting split. This results in a vector of CATE estimates for each individual, $\hat{\theta}_{s}(X)$, $s=1,..,J$. Final CATE estimates are then obtained by taking the median across these estimates, $\widehat{\theta}(X)=median\{\widehat{\theta}_{s}(X)\}$ (Chernozhukov et al., 2018; Jacob, 2020).
@@
Finally, when implementing the mEP-learner, the sieve basis used within the iTMLE process must be specified. The existing iTMLE implementations (Vansteelandt and Morzywołek, 2023; van der Laan et al., 2024) use a univariate trigonometric cosine polynomial basis, as it offers strong approximation guarantees under smoothness assumptions (Zhang and Simon, 2023). We also implement this sieve basis, following the guidance of Zhang and Simon (2023) to define the dimension of the sieve and its default interaction order. However, a cross-validation option for choosing the tuning parameters of this sieve was presented by van der Laan et al. (2024). In addition, we note that when implementing the iTMLE procedure, penalization methods, rather than a linear regression, can be used in the updating step. This was demonstrated by (Vansteelandt and Morzywołek, 2023), and can allow for better model performance when the sieve grows very large.
@@ 3.3.2 Decisions required for assessing CATE performance
The decisions outlined above enable users to obtain CATE estimates from the mDR-learner and mEP-learner. In practice, however, users may also want to assess the accuracy of these estimates by obtaining measures of uncertainty (e.g., confidence intervals) or calculating evaluation metrics such as the mean squared error (MSE).
Calculating measurements of uncertainty for CATE estimates generated using non-parametric estimators is challenging, as the theoretical convergence guarantees required for confidence intervals to be derived are often not met when using non-parametric estimation techniques. However, recent work by Takatsu and Westling (2022); Bonvini et al. (2023) and Ritzwoller and Syrgkanis (2024) offer some solutions. Bonvini et al. (2023) provides options for obtaining confidence intervals when kernel based approaches are used to estimate a univariate, additive or has partial dependence CATE, while Ritzwoller and Syrgkanis (2024) offers a half-sample bootstrapping approach which can provide confidence intervals when kernel based approaches are used to estimate the pseudo-outcome regression. As the work we present in this paper is focused on outlining general frameworks for estimating the CATE in the presence of MAR outcome data, we leave the application of these novel approaches to the user.
We also note how evaluating the performance of CATE estimators in real (as opposed to simulated) data can be problematic, as only one of the two potential outcomes, $Y(0)$ and $Y(1)$, is ever observed for each individual. Consequently, calculating standard metrics such as the MSE in practice is not possible, and instead, estimators of these metrics must be constructed. Simple estimators of the MSE, using either potential outcome predictions on their own, or in combination with IPTWs have been proposed (Schuler et al., 2018), however, these estimators are not robust to errors in the outcome or propensity score models and are prone to plug-in bias. Equally, EIF based estimators which are typically used to address plug-in bias such as this, encounter their own issues, as by definition, the MSE metric and its EIF will be 0 when CATE estimates are unbiased. Alternatively, model performance can be assessed by reviewing the stability of model estimates across different sample splitting seeds, enabling users to understand the variability in individual CATE estimates obtained from estimators such as the mDR-learner or mEP-learner.
@@ 4.1 Set up
We study the finite sample performance of the mDR-learner and mEP-learner across three data generating processes (DGPs), where each DGP corresponds to a setting in which missing outcome data can complicate the estimation of the CATE.
In each setting, we generate 6 uniformly distributed covariates $Z$ (independently), a binary treatment $A$ and a continuous outcome $Y$. In the first two DGPs, we define a simple unexposed outcome function $\mu^{0}(Z)$, a complex CATE $\theta(X)$ and define outcome missingness such that it occurs with high probability in only the treatment arm (DGP 1), or in both arms (DGP 2). This missingness mechanism makes the complex CATE more challenging to learn. In the third DGP, we define a complex unexposed outcome function $\mu^{0}(Z)$, a simple CATE $\theta(X)$ and allow outcome missingness to occur evenly across treatment arms, now making the simple CATE difficult to learn. More details can be found in Appendix G.
We vary the training data sample size from 400 to 3200, and use 500 replicates for each scenario. We compare the mDR-learner and mEP-learner to three alternative CATE estimators; the DR-learner, EP-learner and the T-learner, each implemented using (1) available cases (Section 2.3) and (2) imputed outcomes, obtained from imputation models. All estimators were implemented using 10 fold cross-validation and all models (including imputation models) were fitted using the Super Learner (Van der Laan et al., 2007), which included linear regression, LASSO, random forest and neural nets, amongst other algorithms in its library. Further details, including tuning parameters are reported in Appendix H.
To assess the performance of each estimator, we generated one test dataset with sample size $n=10,000$ per DGP. For each estimator applied to each DGP we obtained the CATE estimates for each individual in this test dataset. Performance was measured by calculating the root mean square median error (RMSME) of each learner (Appendix I). We use this metric as mean root mean square error (RMSE) estimates were found to be skewed when using 500 replications. For comparisons made using the mean RMSE, see Appendix J.
@@ 4.2 Findings
Figure 1 reports the RMSME for each candidate learner. The left column compares the mDR-learner and mEP-learner with the T-learner, DR-learner and EP-learner using available cases. In this comparison we see that the mDR-learner has better performance than the DR-learner across all three DGPs. The same trend can also be seen for the mEP-learner when compared with the EP-learner. The right-hand column shows the performance of the T-learner, DR-learner and EP-learner using imputed outcomes. Here, the mDR-learner and mEP-learner outperformed their corresponding learner with imputed outcomes when the CATE was complex (DGP 1 and 2), demonstrating that the errors introduced by imputing the outcomes propagated through the CATE estimators. However, when the CATE was simple, the DR-learner and EP-learner combined with outcome imputations performed well.
Compared to the available case/imputed outcomes T-learner, the mDR-learner and mEP-learner were less sensitive to specific nuisance function specification, while the T-learner’s performance depends greatly on the complexity of the underlying outcome functions, and worsens substantially when the outcome functions are complex (DGP 3).
@@
Finally, the CATE estimates obtained across simulations for the mEP-learner, EP-learner and T-learner were more stable than those obtained by the mDR-learner and DR-learner. In both the mDR-learner and DR-learner, extreme CATE estimates were produced in some simulated settings, as positivity violations within the folds of the training data occurred. This can be seen more clearly when performance is measured using mean RMSE (Appendix J).
@@
Figure 1: Root mean square median error (RMSME) for mDR-learner, mEP-learner, DR-learner, EP-learner and T-learner in three DGPs plotted by training sample size. Plots in the left column compare the mDR-learner and mEP-learner to the available case versions of the DR-learner, EP-learner and T-learner in DGP 1, 2 and 3 respectively. Plots in the right column compare the mDR-learner and mEP-learner to the DR-learner, EP-learner and T-learner when used in combination with an outcome imputation model in DGP 1, 2 and 3 respectively.
5 ACTG175 trial analysis
5.1 Background
We illustrate the use of the mDR-learner and mEP-learner by applying them to the ACTG175 trial (Hammer et al., 1996), which studied the effect of Zidovudine mono-therapy vs alternative antiretroviral therapies among HIV-1-infected subjects. The ACTG175 trial is a randomised, double blind, placebo controlled trial which recruited 2139 patients from 43 AIDS Clinical Trial Units and 9 National Hemophilia Foundation sites in the United States and Puerto Rico, from 1991 to 1994. Participants were randomised to one of four regimens: Zidovudine mono-therapy; Zidovudine and Didanosine; Zidovudine and Zalcitabine or Didanosine mono-therapy, with the randomisation stratified according to the length of prior antiretroviral therapy.
Baseline covariates were recorded including information on demographics, medical history and disease progression. Lab measurements of CD4 and CD8 T-cell counts (cells/mm3) were taken at baseline as well as at $20\pm 5$ weeks and $96\pm 5$ weeks, with the measurement at $96\pm 5$ weeks defined as the end point of interest. Of the 2139 patients randomised, 797 (37%) were lost to follow-up, either dropping out or dying during the follow up period, however, information on reason for being lost to follow-up is not present in the available data.
We focus on an intention-to-treat comparison between Zidovudine mono-therapy ($n=532$) and all alternative therapy lines ($n=1607$), as done by Wang et al. (2021). We estimate two CATEs, one conditional on all baseline covariates, and one conditional only on age (Fan et al., 2017).
5.2 Methods
CATEs are estimated using the mDR-learner and mEP-learner, and compared to estimates from the DR-learner, EP-learner and T-learner, all fit using available case analyses (Section 2.3). Propensity score, missingness and outcome models, all include the following 12 baseline covariates: age (years), gender (male/female), race (white/non-white), weight (kg), hemophilia (yes/no), homosexual activity (yes/no), Karnofsky score (on a scale of 0-100), symptomatic status (symptomatic/asymptomatic), history of intravenous drug use (yes/no), antiretroviral history (experienced/naive), CD4 T-cell count (cells/mm3) at baseline and CD8 T-cell count (cells/mm3) at baseline. All nuisance models are implemented with a Super Learner and details on the libraries/tuning parameters used within the Super learner are outlined in Appendix H. All methods are implemented using 10 fold cross-validation and we report the median CATE estimate over 50 different sample splitting random seeds.

### v2

@@ 3 DR-learner/EP-learner extensions
In this section, we extend the DR-learner (Kennedy, 2023) and EP-learner (van der Laan et al., 2024) to handle missing outcome data, leading to two new estimators, the mDR-learner (missing outcome DR learner) and mEP-learner (missing outcome EP learner) respectively.
@@
We begin by considering an extension of the DR-learner to the MAR outcome data setting. Recall that the DR-learner does not account for the under-representation that occurs as a result of MAR outcomes. This is because the DR-learner’s pseudo-outcomes are derived using a risk function which assumes complete data. However, when outcomes are MAR, the risk function takes a new form, $\psi_{CATE}=\operatorname{\mathbb{E}}[\theta(X)^{2}-2\theta(X)(\operatorname{\mathbb{E}}[Y|C=1,A=1,Z]-\operatorname{\mathbb{E}}[Y|C=1,A=0,Z])]$, which now involves the indicator for outcomes being non-missing. It can be shown (Appendix D) that the EIF is
@@
Using this EIF, we can construct pseudo-outcomes for the mDR-learner as:
@@
We note that IPCWs now appear in alongside IPTWs, allowing these pseudo-outcomes to account for the shift in covariate distribution caused by both missing outcome data and confounding. The mDR-learner then proceeds by regressing the estimates of these pseudo-outcomes $Y_{mDR}$ against covariates $X$ to obtain estimates of the CATE (See Algorithm 1).
Figure 1: mDR-learner algorithm
@@
For the mDR-learner to obtain oracle efficiency, it requires that two conditions hold; firstly, that the product of the convergence rates for the outcome predictions and propensity score estimates is faster than the rate for the oracle estimator, and secondly, that the product of the convergence rates for the outcome predictions and missingness estimates is also faster than the rate for the oracle estimator. For instance, consider an example where the oracle CATE estimator converges at a $\sqrt{n}$ rate. In this example, if we wish to estimate the nuisance functions within the mDR-learner without impacting the overall convergence rate of our CATE estimator, we require that the estimates of each of the nuisance functions converge at rates faster than $4]{n}$, hence allowing the two products of these rates to converge faster than $\sqrt{n}$. Equally, when the outcome predictions converge at rates slower than this, oracle efficiency can still be obtained as long as the propensity score estimates and missingness function estimates converge at rates which are sufficiently fast for the above conditions to hold. Under these weakened convergence requirements, ML can then be used when estimating nuisance functions within the mDR-learner without errors propagating through to the CATE.
@@
We now demonstrate how the EP-learner, a targeted-learning based framework, can be extended for the setting with MAR outcome data. When outcomes are MAR, the EP-learner fails to account for under-representation introduced by missing outcomes. If we wish to account for this, we must recalculate the EIF which it uses to derive its pseudo-outcomes, with the new EIF taking a new form, eq. (7). This EIF contains IPCWs within its drift term (second term), and hence, if we wish to set the sample average of the drift term to 0, removing plug-in bias, we must update the iTMLE process used within the EP-learner.
To do so, we re-define the weight used in the iTMLE algorithm by considering eq. (7), defining the weight to be $\hat{H}(A,C,Z)=\frac{CA}{\hat{G}(A,Z)\hat{\pi}(Z)}+\frac{C(1-A)}{\hat{G}(A,Z)\left(1-\hat{\pi}(Z)\right)}$. This weight ensures that the sample average of the second term in eq. (7) converges to 0, and hence the mEP-learner will also be oracle efficient when its nuisance function estimates converge sufficiently fast.
Figure 2: mEP-learner algorithm
@@ 3.3 Implementation
As both the mDR-learner and mEP-learner are general frameworks for estimating the CATE with MAR outcome data, their implementation requires the user to make several key decisions. In this section, we break these decisions down into two groups; (a) decisions required for obtaining CATE estimates, and (b) decisions required for assessing CATE performance.
@@ 3.3.1 Decisions required for obtaining CATE estimates
As both learners are model-agnostic the user must first decide how to estimate the nuisance functions/pseudo-outcome model. Data adaptive techniques can be chosen, however, if oracle performance is to be achieved, the estimates from the nuisance models must converge sufficiently fast to the truth. For this reason, we illustrate their implementation using the Super Learner (Van der Laan et al., 2007), an ensemble learner that allows a range of data-adaptive algorithms to be implemented and which performs asymptotically as well as its best candidate learner.
@@
After algorithm choice, users are required to decide which type of sample splitting they will implement within the mDR or mEP learners, ensuring the data used in the pseudo-outcome regression is i.i.d. Various options exist for achieving this, including K-fold cross-fitting (see Section 2.2) or independent sample splitting (Kennedy, 2023), however, estimators that allocate fully independent data for each nuisance/target model are typically less efficient, introducing finite sample bias by reducing samples sizes. For this reason, we demonstrate the mDR-learner and mEP-learner using a 10-fold cross-fitting process (see Algorithm 1), but note that the appropriate number of folds will depend on the complexity and smoothness of the underlying nuisance/target functions. If users wish to explore variations of cross-fitting approaches, sensitivity analyses could be run using an alternative number of splits.
@@
We also highlight that when cross-fitting is used, reduced training sample sizes for each model can introduce positivity violations, leaving certain subgroups poorly represented within some folds. To overcome this, the mDR-learner and mEP-learner can be run multiple times ($J$ times), using a different seed for the cross-fitting split. This results in a vector of CATE estimates for each individual, $\hat{\theta}_{s}(X)$, $s=1,..,J$. Final CATE estimates are then obtained by taking the median across these estimates, $\widehat{\theta}(X)=median\{\widehat{\theta}_{s}(X)\}$ (Jacob, 2020).
@@
Finally, when implementing the mEP-learner, the sieve basis used within the iTMLE process must be specified. The existing iTMLE implementations use a univariate trigonometric cosine polynomial basis, as it offers strong approximation guarantees under smoothness assumptions (Zhang and Simon, 2023). We also implement this sieve basis, following the guidance of Zhang and Simon (2023) to define the dimension of the sieve and its interaction order. However, alternative options exist, such as a cross-validation option for choosing the sieves tuning parameters (van der Laan et al., 2024) and a penalized iTMLE implementation, which can improve performance for large sieves (Vansteelandt and Morzywołek, 2023).
@@ 3.3.2 Decisions required for assessing CATE performance
In addition to obtaining CATE estimates, users may also want to assess the accuracy of these estimates by obtaining measures of uncertainty or calculating evaluation metrics. Calculating measurements of uncertainty for CATE estimates generated using non-parametric estimators is challenging, as the theoretical convergence guarantees required for confidence intervals (CIs) to be derived are often not met. However, recent work by Takatsu and Westling (2022); Bonvini et al. (2023) and Ritzwoller and Syrgkanis (2024) offer potential solutions. In this work we explore how one of these techniques can be used with our estimators, focusing on the half-sample bootstrap approach (Ritzwoller and Syrgkanis, 2024), which can provide CIs when kernel based approaches are used to estimate the pseudo-outcome regression. Further details on this technique are outlined in Appendix G, and we evaluate the performance of this approach within our simulation study (See Section 4) in Appendix L.
Additionally, we note that evaluating the performance of CATE estimators when using real world data is challenging, with only one of the two potential outcomes, $Y(0)$ and $Y(1)$, ever observed for each individual. This means standard metrics such as the MSE cannot be calculated and used for estimator comparison, and equally makes deriving EIF based estimators for these metrics tricky, as by definition, the MSE metric and its EIF will be 0 at the true CATE. Alternatively, model performance can be assessed by reviewing the stability of model estimates across different sample splitting seeds, enabling users to understand the variability in individual CATE estimates. Additionally, if choosing between the mDR/mEP learners, users should focus on the strengths/weaknesses of each learner, with the mDR-learner performing best when the mEP-learner’s sieve poorly approximates the CATE, i.e, when the CATE is non-smooth or sparse, while the mEP-learner may perform best when the mDR-learner’s weights and pseudo-outcomes are highly variable/unstable.
@@ 4.1 Set up
We study the finite sample performance of the mDR-learner and mEP-learner across three data generating processes (DGPs), where each DGP corresponds to a setting in which missing outcome data can complicate the estimation of the CATE. In each setting, we generate 6 uniformly distributed covariates $Z$, a binary treatment $A$ and a continuous outcome $Y$. In the first two DGPs, we define a simple unexposed outcome function $\mu^{0}(Z)$, a complex CATE $\theta(X)$ and define outcome missingness such that it occurs with high probability in only the treatment arm (DGP 1), or in both arms (DGP 2). This makes the complex CATE challenging to learn. In the third DGP, we define a complex unexposed outcome function, a simple CATE and define outcome missingness to occur with high probability in only the treatment arm, making the simple CATE difficult to learn (Appendix H).
We vary the training data sample size from 400 to 3200, and use 500 replicates for each scenario. We compare the mDR-learner and mEP-learner to four alternative CATE estimators; the IPTW-IPCW learner, the DR-learner, EP-learner and the T-learner, with the later three implemented using (1) available cases (Section 2.3) and (2) imputed outcomes. All estimators were implemented using 10 fold cross-fitting, with the nuisance models fit using the Super Learner and the pseudo-outcome models fit using random forests (with 500 bootstrap half samples), enabling the generation of half-sample bootstrap CIs. To assess the performance of each estimator, we generated one test dataset with sample size $n=10,000$ per DGP and obtained the CATE estimates for each individual using each estimator. Performance was measured by calculating the root mean square median error (RMSME) of each learner (Appendix J), as mean root mean square error (RMSE) estimates were found to be skewed when using 500 replications. Conditional CI coverage was calculated/reported in Appendix L. For comparisons made using the mean RMSE, see Appendix K.
@@ 4.2 Findings
When comparing the mDR-learner and mEP-learner with the DR-learner and EP-learner using available cases (Figure 3 - Right column), we see the mDR-learner and mEP-learner outperform the DR/EP learners respectively across all three DGPs. Equally, when comparing the mDR-learner and mEP-learner to the DR-learner and EP-learner fit using imputed outcomes (Figure 3 - Left column) both learners outperformed their corresponding imputed outcome version when the CATE was complex (DGP 1 and 2), while when the CATE was simple, the imputed outcome DR-learner and EP-learner performed well. We also note how the IPTW-IPCW learner and available case/imputed outcome T-learner were sensitive to nuisance function complexity, with their performance depending heavily on the complexity of the outcome functions and propensity score/censoring functions respectively. IPTW-IPCW learner results are excluded from Figure 3 to aid interpretability (See Appendix K).
@@
Finally, we note how the CATE estimates obtained across simulations were more stable for the mEP-learner, EP-learner and T-learner compared to those obtained by the mDR-learner, DR-learner or IPTW-IPCW learner. This demonstrates how these estimators are prone to producing extreme CATE estimates when their weights are unstable. This can be seen more clearly when performance is measured using mean RMSE (Appendix K).
@@
Figure 3: Root mean square median error (RMSME) for mDR-learner, mEP-learner, DR-learner, EP-learner and T-learner in three DGPs plotted by training sample size. Plots in the left column compare the mDR-learner and mEP-learner to the DR-learner, EP-learner and T-learner when used in combination with an outcome imputation model in DGP 1, 2 and 3 respectively. Plots in the right column compare the mDR-learner and mEP-learner to the available case versions of the DR-learner, EP-learner and T-learner in DGP 1, 2 and 3 respectively.
5 GBSG2 trial analysis
5.1 Background and methods
We illustrate the use of the mDR-learner and mEP-learner by applying them to the German breast cancer study (GBSG2) randomized trial (Li et al., 2024). This trial randomly assigned patients to a hormonal therapy (n=440) or no hormonal therapy (n=246) after surgery and recorded baseline covariates on demographics, medical history and disease progression. Treatment efficacy was explored by recording recurrence free survival times, and we define our outcome of interest as breast cancer recurrence or death within three years of surgery. As some patients leave the study before making it to three years, missing outcome data is present, with 158 (46.5%) and 66 (26.8%) of the randomized patients lost to follow-up in each treatment arm. We conduct an intention-to-treat analysis and estimate two CATEs: one conditioned on all baseline covariates and one conditioned solely on progesterone receptor levels (fmol/l), where higher levels are associated with greater benefits from hormonal therapies.
In this trial, patients with non-missing outcomes had higher average progesterone receptor levels at baseline than the full randomized population, with a greater increase seen in amongst patients in the hormonal therapy arm. If left unaccounted for, this may result in CATE estimates which suggest hormonal therapies have a greater benefit than is true. Instead, we estimate CATEs using the mDR-learner and mEP-learner, and compare these to estimates from the the DR-learner, EP-learner, T-learner and IPTW-IPCW learner, with the first three fit using available case analyses as well as in combination with imputed outcomes. All nuisance models were fit using all baseline covariates and all models, including the pseudo-outcome models were estimated using a Super Learner, with the focus on obtaining accurate point estimates rather than CIs. The DR, EP and IPTW-IPCW learners are implemented using 10 fold cross-fitting and we report the median CATE estimate over 10 different sample splitting random seeds. Estimates with CIs that were obtained using untuned random forests (and 500 bootstrap samples) are reported in Appendix M.